# Can Large Language Models Generate High-quality Patent Claims?

### v2

@@ 1 Introduction
Large language models (LLMs) have demonstrated remarkable capabilities across a broad spectrum of general tasks, ranging from natural language understanding to complex problem-solving (Zhao et al., 2023; Min et al., 2023). However, the application of LLMs in specialized domains remains under-explored, such as the patent literature (Jiang and Goetz, 2024). Patents, which are legal documents detailing and protecting inventions to promote technical innovations (Mossoff, 2000), present unique challenges due to their complex and technical nature. Meanwhile, their language features a precision level that is unheard of in other text corpora. The potential for LLMs to revolutionize this field is significant, offering possibilities to enhance technical knowledge extraction, patent analysis, and generation (Jiang and Goetz, 2024).
@@
RQ2: Does current patent-specific LLM outperform general LLMs on claim generation? It is shown that small-scale domain-specific LLMs (e.g., 7B) can outperform strong general LLMs (e.g., GPT-4) on specific tasks, such as in math reasoning (Shao et al., 2024). Meanwhile, the legal-specific LLM SaulLM-7B is reported to outperform Mistral-7B and Llama-2-13B on LegalBench-Instruct and Legal-MMLU benchmarks (Colombo et al., 2024). Therefore, we included legal and patent-specific LLMs to investigate whether domain-specific LLMs can perform better on this task.
@@ 2 Related Work
Despite the potential of LLMs, current research in patent writing is still nascent and generally falls short of expectations. There are previous works that have focused on claim generation. A pioneering study by Lee and Hsiang (2020) investigated the use of GPT-2 (Radford et al., 2019) for generating patent claims, finding that minimal training was sufficient for producing patent-like texts, though without evaluating their quality. Subsequently, Lee (2020) broadened this research by training GPT-2 to transform one component of a patent application into another, such as converting abstracts into claims. As abstracts are normally generic and imprecise, generating claims from abstracts may not be a well-conditioned task.
@@ 4.3 RQ3: How do LLMs perform on patent claims’ completeness, clarity, consistency, and logical linkage?
However, the models’ ability to generate the subsequent dependent claims is much worse. Llama-3 substantially fails to capture the features claimed in the referenced dependent claims. Three main weaknesses are raised in the usage of Llama-3. (1) Llama-3 produces claims that merely reference another claim without specifying additional features, rendering these claims redundant and devoid of any added value. (2) Llama-3 sometimes misclassifies the feature types. For example, Llama-3 tends to misclassify device-specific features as procedural features, using active formulations. This misclassification is a serious flaw as it contradicts standard practice and would likely lead to objections during patent examination. (3) Llama-3 generates some repetitive claims, where exactly the same claims are continuously generated multiple times.
@@
Nonetheless, GPT-4 results in the best human evaluation. The primary claim is mostly accurate, and the dependent claims exhibit better technical relevance, clarity, and logical grouping. The model demonstrates a better understanding of technical features, their interactions, and their significance within the overall context of the invention, indicating its capability to evaluate the entire text and generate a coherent claim set. Due to the large inconsistency between human and existing automatic evaluation methods, a promising research direction is to develop new automated metrics that have closer alignment to human evaluation, such as using LLM-based evaluators (Liu et al., 2023b).
@@ Appendix A Patent Description, Abstract, and Claims
Appendix B Model Details
@@ Appendix D More Results
D.1 Statistics of Generated Claims
D.2 Evaluation on Abstracts
@@
D.3 Evaluation on IPC Labels
Normal Prompt Alternative Prompt
Model BLEU R-1 R-L BERTScore BLEU R-1 R-L BERTScore
SaulLM-7B 12.86 30.68 23.07 80.24 12.99 31.45 23.71 80.41
Mistral-7B 29.70 49.17 36.20 85.33 37.87 59.72 43.49 87.81
Llama-3-8B 35.42 57.17 38.88 88.49 35.61 57.50 39.17 88.51
Llama-3-70B 36.40 59.89 41.79 87.44 37.71 60.43 41.89 87.86
Mixtral-8$\times$7B 37.03 60.18 43.67 88.51 37.86 60.04 43.25 88.53
GPT-4 15.73 52.59 37.73 87.34 28.39 59.53 40.61 87.72
Table 11: Automated evaluation results on different prompts.
D.4 Evaluation on Alternative Prompts

### v3

@@ 1 Introduction
Large language models (LLMs) have demonstrated remarkable capabilities across a broad spectrum of general tasks, ranging from natural language understanding to complex problem-solving (Zhao et al., 2023; Min et al., 2023). However, the application of LLMs in specialized domains remains under-explored, such as the patent literature (Jiang and Goetz, 2025). Patents, which are legal documents detailing and protecting inventions to promote technical innovations (Mossoff, 2000), present unique challenges due to their complex and technical nature. Meanwhile, their language features a precision level that is unheard of in other text corpora. The potential for LLMs to revolutionize this field is significant, offering possibilities to enhance technical knowledge extraction, patent analysis, and generation (Jiang and Goetz, 2025).
@@
Patent-CR (Jiang et al., 2025b) Claim revision Claim $\rightarrow$ claim 22K
@@
RQ2: Does current patent-specific LLM outperform general LLMs on claim generation? It is shown that small-scale domain-specific LLMs (e.g., 7B) can outperform strong general LLMs (e.g., GPT-4) on specific tasks, such as in math reasoning (Shao et al., 2024) and fact-checking (Zhang et al., 2024a). The legal-specific LLM SaulLM-7B is also reported to outperform Mistral-7B and Llama-2-13B on LegalBench-Instruct and Legal-MMLU benchmarks (Colombo et al., 2024). Therefore, we included legal and patent-specific LLMs to investigate whether domain-specific LLMs can perform better on this task.
@@ 2 Related Work
Despite the potential of LLMs, current research in patent writing is still nascent and generally falls short of expectations. Previous works have focused on claim generation. A pioneering study by Lee and Hsiang (2020) investigated the use of GPT-2 (Radford et al., 2019) for generating patent claims, finding that minimal training was sufficient for producing patent-like texts, though without evaluating their quality. Subsequently, Lee (2020) broadened this research by training GPT-2 to transform one component of a patent application into another, such as converting abstracts into claims. As abstracts are normally generic and imprecise, generating claims from abstracts may not be a well-conditioned task. A concurrent work investigated whether LLMs can revise patent claims to improve quality (Jiang et al., 2025b). A follow-up study further explored claim generation on European patents (Jiang et al., 2025a).
@@ 4.3 RQ3: How do LLMs perform on patent claims’ completeness, clarity, consistency, and logical linkage?
However, the model’s ability to generate the subsequent dependent claims is much worse. Llama-3 substantially fails to capture the features claimed in the referenced dependent claims. Three main weaknesses are raised in the usage of Llama-3. (1) Llama-3 produces claims that merely reference another claim without specifying additional features, rendering these claims redundant and devoid of any added value. (2) Llama-3 sometimes misclassifies the feature types. For example, Llama-3 tends to misclassify device-specific features as procedural features, using active formulations. This misclassification is a serious flaw as it contradicts standard practice and would likely lead to objections during patent examination. (3) Llama-3 generates some repetitive claims, where exactly the same claims are continuously generated multiple times.
@@
Nonetheless, GPT-4 results in the best human evaluation. The primary claim is mostly accurate, and the dependent claims exhibit better technical relevance, clarity, and logical grouping. The model demonstrates a better understanding of technical features, their interactions, and their significance within the overall context of the invention, indicating its capability to evaluate the entire text and generate a coherent claim set. Due to the large inconsistency between human and existing automatic evaluation methods, a promising research direction is to develop new automated metrics that have closer alignment to human evaluation, such as using LLM-based evaluators (Liu et al., 2023b; Jiang et al., 2025c).
@@ 5 Conclusion
To enhance the quality of generated claims, future research could explore recent advancements in uncertainty estimation for LLMs (Zhang et al., 2024b, c; Yang et al., 2024). By identifying uncertain generations, human verification can refine the output and improve performance beyond fully automated generation.
@@ Appendix B Model Details
Normal Prompt Alternative Prompt
Model BLEU R-1 R-L BERTScore BLEU R-1 R-L BERTScore
SaulLM-7B 12.86 30.68 23.07 80.24 12.99 31.45 23.71 80.41
Mistral-7B 29.70 49.17 36.20 85.33 37.87 59.72 43.49 87.81
Llama-3-8B 35.42 57.17 38.88 88.49 35.61 57.50 39.17 88.51
Llama-3-70B 36.40 59.89 41.79 87.44 37.71 60.43 41.89 87.86
Mixtral-8$\times$7B 37.03 60.18 43.67 88.51 37.86 60.04 43.25 88.53
GPT-4 15.73 52.59 37.73 87.34 28.39 59.53 40.61 87.72
Table 9: Automated evaluation results on different prompts.
@@ Appendix D More Results
D.1 Statistics of Generated Claims
D.2 Evaluation on Alternative Prompts
@@
D.3 Evaluation on IPC Labels
D.4 Evaluation on Abstracts